The distance from antville to beetletoon is 168km. Mr caterpillar travels from antville to beetletoon at an average speed of 28km/h. How long did it take him?

Answer by rothauserc(4718)

rate * time = distance
time = distance / rate = 168 / 28 = 6 hours
Mr caterpillar took 6 hours to travel to beetletoon from antville
check answer
28 * 6 = 168
168 = 168
our answer checks